OpenStudy (anonymous):

Find 2 vectors that are orthogonal to v=<2,-3>

OpenStudy (anonymous):

Recall that orthogonal vectors will have a dot product of 0. \[\implies <2,-3>\cdot<a,b> = 2a + -3b = 0\] Pick some solutions.

OpenStudy (anonymous):

ya but i got two unknown

OpenStudy (bobbyleary):

Orthogonal essentially means perpendicular. One way to visualize this is by plotting <2,-3>, then drawing a perpendicular to this. One solution is <-3,-2>. To check your answer, the dot product of these vectors must equal zero, so <2,-3>(dot)<-3,-2> = 2*-3 + -3*-2 = 0. polpak will give you a better solution though...:)

OpenStudy (anonymous):

is it possible if i want to get the answer by calculation?

OpenStudy (anonymous):

Having two unknowns just means that there is a whole bunch of solutions.

OpenStudy (anonymous):

\[2a - 3b = 0 \implies 2a = 3b \implies b = \frac{2}{3}a\] Pick any a you like. Make b be 2/3's of that. Boom, orthogonal vector.

OpenStudy (anonymous):

ok ..thx